OpenStudy (tazzwilliams):

F(x) = x 2 + 3x - 2 F(a) = @Qwertty123

OpenStudy (qwertty123):

So do you know how to solve this?

OpenStudy (tazzwilliams):

no not really

OpenStudy (qwertty123):

okay.. So you want to start off with is it x^2 or x 2?

OpenStudy (tazzwilliams):

x2

OpenStudy (qwertty123):

oh.. Simplifying: F(x) = x * 2 + 3x + -2 Multiply F * x xF = x * 2 + 3x + -2 Reorder the terms for easier multiplication: xF = 2x + 3x + -2 Reorder the terms: xF = -2 + 2x + 3x Combine like terms: 2x + 3x = 5x xF = -2 + 5x

OpenStudy (qwertty123):

Solving xF = -2 + 5x Move all terms containing x to the left, all other terms to the right. Add '-5x' to each side of the equation. -5x + xF = -2 + 5x + -5x Combine like terms: 5x + -5x = 0 -5x + xF = -2 + 0 -5x + xF = -2 Reorder the terms: 2 + -5x + xF = -2 + 2 Combine like terms: -2 + 2 = ?

OpenStudy (tazzwilliams):

ok i think i got it now

OpenStudy (tazzwilliams):

thanks